OpenStudy (anonymous):

Given: G = 6.67259 × 10−11 Nm2/kg2 The acceleration of gravity on the surface of a planet of radius R = 7770 km is 12.6 m/s2. What is the period T of a satellite in circu- lar orbit h = 10800.3 km above the surface? Answer in units of s. *note:*(must convert distance to meters)

OpenStudy (anonymous):

\[mg = G\frac{ mM }{R^2 }\]

OpenStudy (anonymous):

\[g = G\frac{ M }{R^2 }\] \[12.6=6.67259 × 10^{−11}\frac{ M }{(7700*1000)^2 }\]

OpenStudy (anonymous):

I dont have mass thats the part that is tripping me up lol

OpenStudy (anonymous):

find M, the mass of the planet, and use in: \[\frac{ V^2 }{ r} = G\frac{ M }{ r^2 }\] (where r is the radius of the orbit) in order to find V. from V, period is easy to find...

OpenStudy (anonymous):

V=\[\sqrt{gr}\]

OpenStudy (anonymous):

then i took that and used T=2piR/V

OpenStudy (anonymous):

then T/3600

OpenStudy (anonymous):

should i have divided by something else to get my period? I:

OpenStudy (anonymous):

doesn't work that way

OpenStudy (anonymous):

the value given for 'g' is only valid on the planet's surface

OpenStudy (anonymous):

it's a crazy fact about gravity that the force of attraction two masses experience is dependent on the inverse square of the distance between them.... if you were on the moon, do you think you would experience an acceleration of 9.8 m/s^2 towards the earth?

OpenStudy (anonymous):

this is universal gravitation

OpenStudy (anonymous):

so g is constant for all space, is that what you're saying? ... cool.

OpenStudy (shamim):

can i try to solve it

OpenStudy (shamim):

actually my equation editor is not working now. its a big problem

OpenStudy (anonymous):

note that \(g=\frac{GM}{r^2}\) g=12.6, G is given, R is 7770000m. I think you can find the M......